A local farmer owns 600 animals. The pie chart below gives the percent of each type owned. How many are pigs?
sheep 15.5%, horses 16.1%, cows 8.7%, chickens 41.2%, pigs 18.5%
please help!

Answer by anomis(18)

18.5/100 * 600 = 18.5 *6 = 111
111 pigs